OpenStudy (anonymous):

Identify the domain of the equation y = x2 − 6x + 1. x ≤ 3 x ≥ −8 x ≥ −2 All real numbers

OpenStudy (anonymous):

lets square it first y=(x2-6x+9)-9+1 y=(x-3)^2-8

jimthompson5910 (jim_thompson5910):

Since you can plug in ANY real number for x, this means that the domain is the set of all real numbers. So the answer is choice D

OpenStudy (anonymous):

all real